# Formal Foundations for Controlled Stochastic Activity Networks

### v1

@@ 3 Nondeterministic Models
The results show that the standard closure properties of $\omega$-regular languages (as established by Büchi [Buchi1962], McNaughton [McNaughton1966], and Landweber [Landweber1969]) remain valid at the language level for controlled Büchi automata. However, uniform syntactic realizations exist only in the regular (deterministic) fragment. This highlights the same hierarchy observed in the finite setting: uniform (model-level) closure becomes strictly weaker than standard (semantic) closure once non-determinism or infinite-memory control is introduced.
@@
Proof Sketch.
For $\Pi\in\{0,F\}$, controlled Büchi automata reduce to classical nondeterministic Büchi automata. The emptiness problem for these automata is PSPACE-complete. It can be decided using graph-search algorithms that detect reachable accepting cycles, such as nested depth-first search (DFS) [baier2008principles, Vardi1985, Buchi1962].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown Büchi automaton. The emptiness problem remains decidable via saturation techniques and automata-theoretic constructions developed for pushdown systems [Esparza1997PDABuchi, Walukiewicz2001, Bouajjani2000]. Decidability assumes a finite stack alphabet and a finite-state plant; the Büchi set is evaluated on the standard product pushdown system.
For $\Pi\in\{\textit{Tape},H\}$, the control mechanism can simulate a Turing machine on infinite inputs, and the emptiness problem becomes equivalent to the nonemptiness problem for $\omega$-languages accepted by Turing machines, which is undecidable [Sipser2013, Hopcroft1979]. ∎
@@ 4 Probabilistic Models
We now present a notion of equivalence for controlled probabilistic automata which is similar to the one proposed for probabilistic automata [Larsen1991].
@@
  • $\gamma:M\to Dist(C)$: probabilistic control function.
@@
  • $\gamma:\Gamma\to Dist(C)$: probabilistic control output function depending only on the top stack symbol.
@@
  • $\gamma:\Gamma\to Dist(C)$: probabilistic control output function depending only on the current scanned tape symbol.
@@
where $\Pr_{\pi}(\rho)$ is defined as in Definition 4.13 above.
@@
$L^{\theta}_{0}(U)\subset L^{\theta}_{F}(U)\subset L^{\theta}_{\text{Stack}}(U)\subset L^{\theta}_{\text{Tape}}(U)\subset L^{\theta}_{H}(U),$
$L^{\theta}_{0,P}(U)\subset L^{\theta}_{F,P}(U)\subset L^{\theta}_{\text{Stack},P}(U)\subset L^{\theta}_{\text{Tape},P}(U)\subset L^{\theta}_{H,P}(U).$
@@
Proof Sketch.
Analogous to the proof sketch of Theorem 3.29. ∎
Theorem 4.20.
Proof Sketch.
@@
Proof Sketch.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was first established by Rabin [Rabin1963] through a reduction from the Post correspondence problem. Paz [Paz1971] provides a comprehensive exposition of this result and related decision problems, including the structure of probabilistic languages and the behavior of cutpoints. Madani, Hanks, and Condon [madani2003] later extended these ideas to probabilistic planning, showing that undecidability persists in control-oriented settings. Together, these results establish that no algorithm can decide emptiness for probabilistic automata under a non-strict cutpoint semantics.
@@
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic Büchi automaton, and let $\pi:(Q\times A)^{*}\to\text{Dist}(C)$ be a probabilistic policy. We can construct a probability space $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$ under policiy $\pi$ as follows. Define the sample space of infinite runs as $\Omega=(Q\times A)^{\omega}$, and let $\mathcal{F}$ be the $\sigma$-algebra generated by cylinder sets [billingsley1995probability] of the form:
@@
where $h_{i}=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{i-1},a_{i-1})$ is the prefix history leading up to time step $i$. By Kolmogorov’s extension theorem [Kolmogorov1933, billingsley1995probability], this defines a unique probability measure $\mathbb{P}_{\pi}$ on $(\Omega,\mathcal{F})$.
@@
where $\text{Inf}(\rho)$ is the set of states that appear infinitely often in the projection of $\rho$ onto $Q$. Given a threshold $\theta\in[0,1]$, the probabilistic $\omega$-language accepted under a probabilistic policy $\pi$ by $U_{\omega}$ is defined as:
@@
$L^{\omega,\theta}_{0}(U_{\omega})\subset L^{\omega,\theta}_{F}(U_{\omega})\subset L^{\omega,\theta}_{\text{Stack}}(U_{\omega}\subset L^{\omega,\theta}_{\text{Tape}}(U_{\omega})\subset L^{\omega,\theta}_{H}(U_{\omega}),$
$L^{\omega,\theta}_{0,P}(U_{\omega})\subset L^{\omega,\theta}_{F,P}(U_{\omega})\subset L^{\omega,\theta}_{\text{Stack,P}}(U_{\omega})\subset L^{\omega,\theta}_{\text{Tape,P}}(U_{\omega})\subset L^{\omega,\theta}_{H,P}(U_{\omega}).$
@@
When $\theta=0^{+}$ or $\theta=1$, where $0^{+}$ denotes a sufficiently small positive real number, the model is said to have a positive acceptance or almost-sure acceptance semantics, respectively [baier2012probabilistic].
We now consider the global language families associated with controlled probabilistic Büchi automata as formalized in Definition  4.28. We define sets of language classes indexed by controlled probabilistic systems and stratified by policy type.
@@
Let $\mathcal{U}_{\omega}$ be the set of all probabilistic controlled Büchi automata (Definition  4.26). For each policy class $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
@@
Proof Sketch.
Analogous to the proof sketch of Theorem 3.29. ∎
Proof Sketch.
The argument parallels Theorem 3.29, now lifted to the probabilistic Büchi setting. Each inclusion follows because stronger controller classes strictly extend weaker ones: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-based $\subset$ history-dependent. Strictness is witnessed by adapting the corresponding deterministic constructions of Theorem 3.29, which remains valid under degenerate probabilistic transitions (probability 0 or $1$). Thus, each containment in the stated chain is proper. ∎
@@
Proof Sketch.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was established by Rabin [Rabin1963] for probabilistic finite automata. This result extends to probabilistic Büchi automata (PBAs) under positive semantics, as shown in [baier2012probabilistic, courcoubetis1995complexity]. Specifically, given a PBA $A=(Q,A,\delta,q_{0},F)$ and a threshold $\theta\in(0,1]$, deciding whether there exists $w\in A^{\omega}$ with $\Pr(A\text{ accepts }w)\geq\theta$ is undecidable, even for $\theta=0$.
@@
Since the latter problem is undecidable for PBAs under positive semantics, the emptiness problem for CPBAs is also undecidable. The result continues to hold for deterministic history-dependent policies and fixed thresholds. ∎
@@
The undecidability of the emptiness problem stated above holds for positive semantics, where a word is accepted if its acceptance probability is strictly greater than a threshold $\theta$. In contrast, under almost-sure semantics, where acceptance requires probability $1$, several decision problems become decidable for important subclasses of probabilistic automata [baier2012probabilistic, courcoubetis1995complexity]. This sharp contrast highlights the subtle role of cutpoints in probabilistic verification.
These results highlight the sharp contrast between deterministic and probabilistic settings: while emptiness is tractable for regular and pushdown systems, the introduction of probabilistic thresholds leads quickly to undecidability.
To analyze decision-making in controlled probabilistic systems, it is essential to extend the structural semantics of controlled probabilistic automata with quantitative measures. The reward model introduced below provides such a foundation by associating costs or benefits with both control decisions and activity completions. This framework allows us to formalize performance and optimality criteria, preparing the ground for policy synthesis and complexity analyses in the discrete-time and continuous-time settings explored later in this section.
@@
Let $U=(Q,A,C,P,Q_{0})$ and $D=(Q,C,Q_{0},\lambda,P^{\prime})$ be a controlled probabilistic automaton (CPA) and a discrete-time Markov decision process (DTMDP), respectively, as defined in Definition  4.40 above. Then, there exist policies $\pi_{U}$ and $\pi_{D}$ in $U$ and $D$, respectively, such that the state processes of $U$ and $D$ under policies $\pi_{U}$ and $\pi_{D}$ will be isomorphic respectively.
@@
Proof Sketch.
The existence of optimal stationary deterministic policies for expected total and discounted reward objectives follows from classical results on finite-state MDPs [Howard71, puterman1994markov]. Intuitively, dynamic programming or linear programming methods can compute an optimal policy in polynomial time since the state and action spaces are finite. For time-bounded reachability or reward objectives, PSPACE-completeness follows from standard reductions used in model checking of MDPs [courcoubetis1995complexity], and history-dependent policies may be required to achieve optimality. ∎
@@
For time-bounded objectives, optimality may require history-dependent policies because the choice of action can depend on the remaining time horizon, which cannot be encoded by a stationary deterministic policy alone [courcoubetis1995complexity, baier2008principles].
@@
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [Howard71, puterman1994markov]
Discounted Reward Memoryless deterministic Polynomial time Value iteration / policy iteration [Howard71, puterman1994markov]
Finite-Horizon Reachability / Reward May require history dependence PSPACE-complete Dynamic programming over horizon [courcoubetis1995complexity]
@@
Proof Sketch.
These results follow from the classical theory of Markov decision processes over countable and Borel state spaces. Bounded rewards and non-explosiveness guarantee that the Bellman optimality operator is well-defined and that the value function is finite. For discounted reward criteria, the Bellman operator is a contraction mapping, so convergence of value iteration follows from the Banach fixed-point theorem. Existence of optimal stationary deterministic policies under continuity and compactness assumptions is established using measurable selection theorems. Approximation schemes for infinite-state MDPs are standard techniques for constructing $\varepsilon$-optimal policies. For details, see [puterman1994markov, feinberg2002handbook, HernandezLerma1996]. ∎
@@
Finite-State, Expected Total Reward Memoryless deterministic Polynomial time Linear programming [puterman1994markov]
Finite-State, Discounted Reward Memoryless deterministic Polynomial time Value or policy iteration [puterman1994markov]
Finite-State, Time-Bounded Objectives May require history dependence PSPACE-complete Dynamic programming over horizon [courcoubetis1995complexity]
Infinite-State, Expected or Discounted Reward Stationary deterministic (under continuity) Existence guaranteed; exact computation typically infeasible Value iteration with truncation, uniformization, or grid-based approximation [puterman1994markov]
Infinite-State, Time-Bounded Objectives Often history-dependent or approximate only No general complexity bounds; $\varepsilon$-optimal policies achievable State abstraction and approximation [courcoubetis1995complexity, feinberg2002handbook, HernandezLerma1996]

### v2

@@ 3 Nondeterministic Models
The results show that the standard closure properties of $\omega$-regular languages (as established by Büchi [8], McNaughton [43], and Landweber [39]) remain valid at the language level for controlled Büchi automata. However, uniform syntactic realizations exist only in the regular (deterministic) fragment. This highlights the same hierarchy observed in the finite setting: uniform (model-level) closure becomes strictly weaker than standard (semantic) closure once non-determinism or infinite-memory control is introduced.
@@
Determinization under Restricted Policy Types for $\omega$-Languages
We briefly discuss the extent to which the determinization observation from the finite-word setting carries over to the case of $\omega$-languages.
Theorem 3.70.
Let $S$ be a finite controlled automaton over infinite words, equipped with an $\omega$-regular acceptance condition (e.g., Büchi, parity, Rabin, or Muller), and let the policy type be memoryless (type 0). Then there exists a deterministic controlled automaton $S_{\det}$ over the same input alphabet such that
$\{\,L^{\omega}_{\pi}(S):\pi\in\Pi_{0}(S)\,\}\;=\;\{\,L^{\omega}_{\pi^{\prime}}(S_{\det}):\pi^{\prime}\in\Pi_{0}(S_{\det})\,\}.$
In particular, for type 0 we may, without loss of closed-loop expressiveness over $\omega$-words, restrict attention to deterministic plants.
Proof.
For a fixed plant $S$, the set of memoryless policies $\Pi_{0}(S)$ is finite. Each policy $\pi\in\Pi_{0}(S)$ produces a closed-loop system “$S$ under $\pi$”, which is a finite-state $\omega$-automaton over the input alphabet. By standard results on $\omega$-automata, the induced language $L^{\omega}_{\pi}(S)$ is $\omega$-regular, and there exists a deterministic $\omega$-automaton (for instance, a deterministic parity automaton) $P_{\pi}$ recognizing $L^{\omega}_{\pi}(S)$.
We now combine these finitely many deterministic $\omega$-automata into a single deterministic plant $S_{\det}$ whose control alphabet contains one distinguished action for each policy $\pi\in\Pi_{0}(S)$, intuitively selecting the corresponding component $P_{\pi}$. A memoryless policy on $S_{\det}$ chooses one such action in the initial state and has no further essential influence on the run. By construction, for each $\pi\in\Pi_{0}(S)$ there is a corresponding memoryless policy $\pi^{\prime}\in\Pi_{0}(S_{\det})$ such that $L^{\omega}_{\pi}(S)=L^{\omega}_{\pi^{\prime}}(S_{\det})$, and conversely every closed-loop language of $S_{\det}$ under a memoryless policy arises in this way. This yields the claimed equality of families of $\omega$-languages. ∎
Remark 3.71.
For the finite-memory policy type (type $F$), the situation over $\omega$-languages is more delicate. On a fixed plant $S$, the class of admissible finite-memory controllers may be infinite, so the above finitary construction (which enumerates all policies) does not directly apply. While each individual closed-loop language $L^{\omega}_{\pi}(S)$ is still $\omega$-regular and can be recognized by a deterministic $\omega$-automaton, it is not clear in general whether there always exists a single deterministic controlled automaton $S_{\det}$ whose finite-memory policies generate the same family of closed-loop $\omega$-languages as the original plant under all finite-memory policies. Clarifying this question for type $F$ appears to be an interesting open problem and is left for future investigation.
@@
Proof.
For $\Pi\in\{0,F\}$, controlled Büchi automata reduce to classical nondeterministic Büchi automata. The emptiness problem for these automata is PSPACE-complete. It can be decided using graph-search algorithms that detect reachable accepting cycles, such as nested depth-first search (DFS) [3, 73, 8].
For $\Pi=\textit{Stack}$, the model corresponds to a pushdown Büchi automaton. The emptiness problem remains decidable via saturation techniques and automata-theoretic constructions developed for pushdown systems [16, 74, 6]. Decidability assumes a finite stack alphabet and a finite-state plant; the Büchi set is evaluated on the standard product pushdown system.
For $\Pi\in\{\textit{Tape},H\}$, the control mechanism can simulate a Turing machine on infinite inputs, and the emptiness problem becomes equivalent to the nonemptiness problem for $\omega$-languages accepted by Turing machines, which is undecidable [65, 30]. ∎
@@ 4 Probabilistic Models
We now present a notion of equivalence for controlled probabilistic automata, which is similar to the one proposed for probabilistic automata [37].
@@
  • $\gamma:Q\times M\to Dist(C)$: probabilistic control function.
@@
  • $\gamma:Q\times\Gamma\to Dist(C)$: probabilistic control output function depending only on the top stack symbol.
@@
  • $\gamma:Q\times\Gamma\to Dist(C)$: probabilistic control output function depending only on the current scanned tape symbol.
@@
where $\Pr_{\pi}(\rho)$ is defined as in Definition 4.13 above. When $\theta=0^{+}$ or $\theta=1$, where $0^{+}$ denotes a sufficiently small positive real number, the model is said to have a positive acceptance or almost-sure acceptance semantics, respectively [3].
@@
$L^{\theta}_{0}(U)\subseteq L^{\theta}_{F}(U)\subseteq L^{\theta}_{\text{Stack}}(U)\subseteq L^{\theta}_{\text{Tape}}(U)\subseteq L^{\theta}_{H}(U),$
$L^{\theta}_{0,P}(U)\subseteq L^{\theta}_{F,P}(U)\subseteq L^{\theta}_{\text{Stack},P}(U)\subseteq L^{\theta}_{\text{Tape},P}(U)\subseteq L^{\theta}_{H,P}(U).$
@@
Proof.
@@
Proof.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was first established by Rabin [57] through a reduction from the Post correspondence problem. Paz [53] provides a comprehensive exposition of this result and related decision problems, including the structure of probabilistic languages and the behavior of cutpoints. Madani, Hanks, and Condon [41] later extended these ideas to probabilistic planning, showing that undecidability persists in control-oriented settings. Together, these results establish that no algorithm can decide emptiness for probabilistic automata under a non-strict cutpoint semantics.
@@
Let $U_{\omega}=(Q,A,C,P,Q_{0},F)$ be a controlled probabilistic Büchi automaton, and let $\pi:(Q\times A)^{*}\to\text{Dist}(C)$ be a probabilistic policy. We can construct a probability space $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$ under policiy $\pi$ as follows. Define the sample space of infinite runs as $\Omega=(Q\times A)^{\omega}$, and let $\mathcal{F}$ be the $\sigma$-algebra generated by cylinder sets [5] of the form:
@@
where $h_{i}=(q_{0},a_{0})(q_{1},a_{1})\cdots(q_{i-1},a_{i-1})$ is the prefix history leading up to time step $i$. By Kolmogorov’s extension theorem [34, 5], this defines a unique probability measure $\mathbb{P}_{\pi}$ on $(\Omega,\mathcal{F})$.
@@
where $\text{Inf}(\rho)$ is the set of states that appear infinitely often in the projection of $\rho$ onto $Q$. Given a threshold $\theta\in(0,1]$, the probabilistic $\omega$-language accepted under a probabilistic policy $\pi$ by $U_{\omega}$ is defined as:
@@
When $\theta=0^{+}$ or $\theta=1$, where $0^{+}$ denotes a sufficiently small positive real number, the model is said to have a positive acceptance or almost-sure acceptance semantics, respectively [2].
@@
$L^{\omega,\theta}_{0}(U_{\omega})\subseteq L^{\omega,\theta}_{F}(U_{\omega})\subseteq L^{\omega,\theta}_{\text{Stack}}(U_{\omega}\subseteq L^{\omega,\theta}_{\text{Tape}}(U_{\omega})\subset L^{\omega,\theta}_{H}(U_{\omega}),$
$L^{\omega,\theta}_{0,P}(U_{\omega})\subseteq L^{\omega,\theta}_{F,P}(U_{\omega})\subseteq L^{\omega,\theta}_{\text{Stack,P}}(U_{\omega})\subseteq L^{\omega,\theta}_{\text{Tape,P}}(U_{\omega})\subset L^{\omega,\theta}_{H,P}(U_{\omega}).$
@@
We now consider the global language families associated with controlled probabilistic Büchi automata as formalized in Definition  4.27. We define sets of language classes indexed by controlled probabilistic systems and stratified by policy type.
@@
Let $\mathcal{U}_{\omega}$ be the set of all probabilistic controlled Büchi automata (Definition  4.25). For each policy class $\Pi\in\{0,F,\text{Stack},\text{Tape},H\}$, define:
@@
Proof.
The argument parallels Theorem 3.30, now lifted to the probabilistic Büchi setting. Each inclusion follows because stronger controller classes strictly extend weaker ones: memoryless $\subset$ finite-memory $\subset$ stack-augmented $\subset$ tape-based $\subset$ history-dependent. Strictness is witnessed by adapting the corresponding deterministic constructions of Theorem 3.30 using Lemma 4.33, which remains valid under degenerate probabilistic transitions (probability 0 or $1$). Thus, each containment in the stated chain is proper. ∎
@@
Proof.
The undecidability of the emptiness problem for probabilistic automata with a non-strict cutpoint was established by Rabin [57] for probabilistic finite automata. This result extends to probabilistic Büchi automata (PBAs) as shown in [2, 12]. Specifically, given a PBA $A=(Q,A,\delta,q_{0},F)$ and a threshold $\theta\in(0,1]$, deciding whether there exists $w\in A^{\omega}$ with $\Pr(A\text{ accepts }w)\geq\theta$ is undecidable.
@@
Since the latter problem is undecidable for PBAs, the emptiness problem for CPBAs is also undecidable. The result continues to hold for history-dependent policies and fixed thresholds. ∎
@@
The undecidability of the emptiness problem stated above holds for positive semantics, where a word is accepted if its acceptance probability is strictly greater than zero. In contrast, under almost-sure semantics, where acceptance requires probability $1$, several decision problems become decidable for important subclasses of probabilistic automata [2, 12]. This sharp contrast highlights the subtle role of cutpoints in probabilistic verification.
These results also highlight the sharp contrast between deterministic and probabilistic settings: while emptiness is tractable for regular and pushdown systems, the introduction of probabilistic thresholds leads quickly to undecidability.
To analyze decision-making in controlled probabilistic systems, it is essential to extend the structural semantics of controlled probabilistic automata with quantitative measures. The reward model introduced below provides such a foundation by associating costs or benefits with both control decisions and activity completions. This framework allows us to formalize performance and optimality criteria, preparing the ground for policy synthesis and complexity analyses in the discrete-time and continuous-time settings explored later in this paper.
@@
Let $U=(Q,A,C,P,Q_{0})$ and $D=(Q,C,Q_{0},\lambda,P^{\prime})$ be a controlled probabilistic automaton (CPA) and a discrete-time Markov decision process (DTMDP), respectively, as defined in Definition  4.39 above. Then, there exist policies $\pi_{U}$ and $\pi_{D}$ in $U$ and $D$, respectively, such that the state processes of $U$ and $D$ under policies $\pi_{U}$ and $\pi_{D}$ will be isomorphic respectively.
@@
Proof.
Items (1) and (2) follow from classical results on finite-state discrete-time Markov decision processes. For both expected total reward and discounted reward objectives, there always exists an optimal stationary deterministic policy. Moreover, such a policy can be computed in polynomial time using dynamic programming or linear programming techniques, since the state and action spaces are finite [31, 56].
We now consider time-bounded reachability or reward objectives, i.e., finite-horizon MDPs. When the time bound is a fixed constant or given explicitly, optimal values can be computed efficiently by backward induction over the horizon. However, if the time bound is part of the input and encoded in binary, the numeric value of the horizon may be exponentially large relative to the size of the input. In such cases, even though backward induction remains conceptually valid, evaluating the full dynamic programming recursion may require exponential time.
In this general setting, the problem of computing the optimal value of a finite-horizon MDP can be solved in PSPACE by evaluating the dynamic programming recursion using polynomial space. PSPACE-hardness is established by reductions from space-bounded alternating Turing machines, showing that finite-horizon MDP optimization problems are PSPACE-complete in general [51]. Therefore, time-bounded reachability and reward optimization in finite-state discrete-time MDPs are PSPACE-complete.
Finally, in the time-bounded case, optimal policies may require history dependence, since the choice of an optimal action may depend on the remaining time horizon. Consequently, stationary policies are not sufficient in general for time-bounded objectives, and non-stationary (history-dependent) policies may be required to achieve optimality [56]. ∎
@@
For time-bounded objectives, optimality may require history-dependent policies because the choice of action can depend on the number of remaining steps or time units until the deadline. This dependence introduces a need to track time explicitly, which cannot be captured by stationary policies. In particular, when the optimal decision at a given state varies based on how close the system is to the time bound, a stationary policy—being time-invariant—cannot achieve optimality. Therefore, history-dependent (or time-aware) policies are necessary in general for time-bounded reachability or reward maximization problems.
@@
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [31, 56]
Discounted Reward Memoryless deterministic Polynomial time Value iteration / policy iteration [31, 56]
Finite-Horizon Reachability / Reward May require history dependence PSPACE-complete Dynamic programming over horizon [51]
@@
Proof.
These results follow from the classical theory of Markov decision processes over countable and Borel state spaces. Bounded rewards and non-explosiveness guarantee that the Bellman optimality operator is well-defined and that the value function is finite. For discounted reward criteria, the Bellman operator is a contraction mapping, so convergence of value iteration follows from the Banach fixed-point theorem.
Boundedness of the reward function ensures that the Bellman operator maps bounded functions to bounded functions. Under the assumption of non-explosive dynamics, expected rewards are well-defined. In the case of discounted rewards, the Bellman operator is a $\gamma$-contraction under the supremum norm, and thus value iteration converges to the unique fixed point $V^{*}$, the optimal value function. Approximation techniques for $\varepsilon$-optimal policies include finite-state truncation and quantized abstraction methods, which converge under standard continuity assumptions.
Existence of optimal stationary deterministic policies under continuity and compactness assumptions is established using measurable selection theorems. Approximation schemes for infinite-state MDPs are standard techniques for constructing $\varepsilon$-optimal policies. For details, see [56, 20, 26]. ∎
@@
Finite-State, Expected Total Reward Memoryless deterministic Polynomial time Linear programming [56]
Finite-State, Discounted Reward Memoryless deterministic Polynomial time Value or policy iteration [56]
Finite-State, Time-Bounded Objectives May require history dependence PSPACE-complete Dynamic programming over horizon [51]
Infinite-State, Expected or Discounted Reward Stationary deterministic (under continuity) Existence guaranteed; exact computation typically infeasible Value iteration with truncation, uniformization, or grid-based approximation [56]
Infinite-State, Time-Bounded Objectives Often history-dependent or approximate only No general complexity bounds; $\varepsilon$-optimal policies achievable State abstraction and approximation [51, 20, 26]